1. A C level confidence interval is described as what? (3 points) Oan interval computed from a data sample by a method that has the probability C of producing an interval containing the true value of the parameter any interval within the margin of error O an interval with a margin of error C that is also correct C% of the time Oany interval that follows the 68-95-99.7 rules Oan interval computed from a data sample that guarantees that the probability the interval computed contains the parameter C of nterest 2. A sample of 25 seniors from a large urban school district has a mean Verbal SAT score of 450. (3 points) Assume you know that the standard deviation of the population of Verbal SAT scores for seniors in the district is s-100, A 90% confidence interval for the mean score μ for the population of seniors is used. Which of the following will produce a confidence interval with a smaller margin of error? O Using a confidence interval of 95% OUsing a sample of 100 0 Using an upper critical value of 99% OUsing a sample of 15 seniors OUsing any sample size smaller than 25 3. A sample of 100 janitors in a certain school district found that the average time each janitor (3 points) had worked for the school district was 8 years. Suppose you know that the standard deviation of the population is s 5 years. Construct a 95% confidence interval for the mean time of the population of janitors have spent with the school district 08士1 .98 08 ± 0.82 08 ±0.71 O8 1.53 O8+0.984. Look at the same sample of 25 seniors from the urban school district with the mean and 3 points) standard deviation N(450, 100). A 95% confidence interval for μ for the population of seniors with a margin of error of ± 25 is used. What is the smallest sample size you can take to achieve this same margin of error? 026 045 O50 062 069 Essay Your teacher will grade your response to ensure you receive proper credit for your answer 5, Describe in detail how you would construct a 95% confidence interval for a set of 30 data (4 points) points whose mean is 20 and population standard deviation is 3. Be sure to show that you know the formula and how to plug into it. Also, clearly state the margin of error 6. If the margin of error in an estimate for the mean weight of a shipment is + or-2 pounds at a (4 points) confidence level of 95%, what will be the margin of error at a confidence level of 98%? Be sure to show how you arrived at your answer.
Answer #1

1)
option A)

2)
option B) sample of 100

as sample size increases, margin of error decreases

3)
mean = 8
margin of error = z *sd/sqrt(n)
= 1.96* 5/sqrt(100)
= 0.98

hence
option E) 8 +-0.98 is correct

4)
n= ( z *sd/e)^2
= (1.96 * 100/25)^2
= 61.46
=62

option D)

Please rate

Please post rest question again